### v1

# Disclosure Avoidance for the 2020 Census Demographic and Housing Characteristics File

Abstract.
[14] describe the concepts and methods used by the Disclosure Avoidance System (DAS) to produce formally private output in support of the 2020 Census data product releases, with a particular focus on the DAS implementation that was used to create the 2020 Census Redistricting Data (P.L. 94-171) Summary File. In this paper we describe the updates to the DAS that were required to release the Demographic and Housing Characteristics (DHC) File, which provides more granular tables than other data products, such as the Redistricting Data Summary File. We also describe the final configuration parameters used for the production DHC DAS implementation, as well as subsequent experimental data products to facilitate development of tools that provide confidence intervals for confidential 2020 Census tabulations.
Ryan Cumings-Menon, Robert Ashmead, Daniel Kifer,†$\diamond$
Philip Leclerc, Matthew Spence, Pavel Zhuravlev, John M. Abowd
U.S. Census Bureau
Cornell University $\diamond$ Penn State University
*The views expressed in this technical paper are those of the authors and not those of the U.S. Census Bureau. The Census Bureau has reviewed this data product to ensure appropriate access, use, and disclosure avoidance protection of the confidential source data (Project No. P-7502798, Disclosure Review Board (DRB) approval number: CBDRB-FY23-DSEP-001).
Keywords: Differential Privacy, 2020 Census, TopDown Algorithm, Demographic and Housing Characteristics File
This article is © 2023 by author(s) as listed above. The article is licensed under a Creative Commons Attribution (CC BY 4.0) International license (https://creativecommons.org/licenses/by/4.0/legalcode), except where otherwise indicated with respect to particular material included in the article. The article should be attributed to the author(s) identified above.
1. Introduction
The Census Bureau’s 2020 Disclosure Avoidance System (DAS) was used to produce a formally private file in support of the 2020 Census Redistricting Data (P.L. 94-171) Summary File (hereafter redistricting data); the redistricting data file use case is described in more detail by [14]. The DAS takes the confidential Census Edited File (CEF) as input, which is a set of microdata files with components that include data on individuals, housing units, group quarters (GQs), and households; the output of the DAS is called the Microdata Detailed File (MDF). In the interest of a concise exposition, we refer readers to [14] for more detail on the DAS implementation as well as the concepts in the differential privacy (DP) literature [22] that are used by the DAS. This paper is a companion paper to [14]; we build on this earlier work by providing the updates to the DAS that were required to produce formally private output in support of the Demographic and Housing Characteristics (DHC) File.
The DHC release is an important source of decennial census data for governments, researchers, and businesses because it provides much more granular counts than other sources, such as the redistricting data. This additional granularity supports different use cases than the redistricting data and also raises some technical issues for executing the DAS; the updates described in this paper were motivated by both of these factors. The DHC release is split into two data universes, and counts for each of these two universes are found in distinct DAS implementations. First, the household universe, or DHCH, provides both counts for categories of householders as well as counts for various categories of housing unit and group quarters types. For example, for housing unit types, this includes counts of owner-occupied, vacant, and rented housing units, and, for group quarter types, this includes counts of nursing facilities, college/university student housing, and federal prisons. Second, the persons universe, or DHCP, provides counts of individuals within population groups.
After describing the DAS at a high level in the next subsection, the remainder of the paper is organized as follows. Section 2 describes the schemas and constraints used within DHCH and DHCP DAS implementations. Section 3 describes updates to the computational environment used for the DAS DHC implementation. Section 4 describes both how the DAS team went about supporting the choice of production DHC DAS configuration parameters, as well as the production privacy-loss budget (PLB) allocations used to define the noisy measurement files (NMF) for both types of DHC implementations. Section 5 provides summary metrics for the production 2020 DHC file release. One significant advantage of the formally private methods used by the DAS is that they are sufficiently transparent to allow for accurate inferences, i.e., confidence intervals and hypothesis tests, about the underlying CEF-based tabulations. Section 6 concludes with a description of upcoming experimental data products to support the creation of these inferences.
1.1. A Brief Overview of DAS
Throughout this paper we will make frequent references to [14], but, to make this paper more self-contained, this section outlines the basic steps taken during the DHC and redistricting data. First, after reading the microdata contained in the CEF, these implementations use these confidential microdata to create a flattened fully saturated contingency table, i.e., a flattened histogram, for each census block. In other words, each element (or cell) in the histogram of a given census block corresponds to a possible value of the records, as defined by the product schema, and the value of the cell provides the number of records with this value.
After these histograms have been created in each block geounit, a process called spine optimization is carried out. There are many details related to this process that we will not cover here but that are described by [19]. For the purposes of this paper, it suffices to know that this process defines a hierarchical set of geographic units, which is known as a geographic spine, or simply a spine. For example, the geographic levels for DHC runs, from least to most granular are: US, state, county, prim, tract subset group (TSG), tract subset (TS), optimized block group (OBG), and census block geographic levels. For more detail on the definitions of these geographic levels, see [19]. [^1] State includes state equivalents, namely the District of Columbia. County includes county equivalents. Throughout the paper the geographic unit “block” always means “census block.” The geographic levels used in DHC Puerto Rico runs are the same as the geographic levels listed here, with the exception that a US geographic level is not included. The optimized spine is only used internally within the DAS implementation; the published tabulations do not depend on the definition used for the optimized spine.
After the optimized spine is created, the histogram for each geographic unit (e.g., Fairfax County, Virginia) in the optimized spine is created by summing up the histograms of the blocks it contains. The DAS noisy measurements module, a formally private mechanism that satisfies $\rho-$zero-concentrated differential privacy ($\rho-$zCDP) [16], then forms various marginals from these resulting histograms and adds zero-mean discrete Gaussian noise [17] to these marginals. These noisy marginals are called noisy measurements.
The DAS postprocessing module uses the noisy measurements to create the MDF. Starting from the least granular geographic unit, i.e., the root geographic units, which are either the US or Puerto Rico, it fits a nonnegative and integer histogram to the noisy measurements computed at that geographic level by solving a series of optimization problems, which are described in more detail by [14]. The DAS postprocessing module then fits nonnegative integer histograms to the child geographic units using their noisy measurements, subject to the constraint that their fitted histograms sum to the fitted histogram of the parent geounit. This process is then repeated to define the histogram estimates for each of the remaining geographic levels, from the least to the most granular geographic units. Defining these estimates as solutions to optimization problems also provides a straightforward method to ensure these estimators satisfy a set of constraints, as described by [14]. The specific optimization problems that are used to define the histogram estimate of each geographic unit consist of both least squares and mixed integer programming (MIP) optimization problems. The final step within the DAS implementation is to convert the block geographic level fitted histograms to microdata, which is the output MDF of the DAS implementation.
2. DHC Schemas and Constraints
In the following subsections we begin by describing the schemas and the constraints of both DHCH and DHCP. Both of these schemas are more granular than their respective redistricting data counterparts. We also ensure that the DHCP MDF is consistent with the MDF of the persons redistricting DAS output and that the DHCH MDF is consistent with the MDF of the units redistricting DAS output. In other words, it is possible to recreate the redistricting persons universe MDF using only the DHCP MDF and to recreate the redistricting units universe MDF using only the DHCH MDF.
2.1. DHCH Schema
Before providing the schema for the DHCH histogram, we define the DHCH attribute HHTYPE, i.e., household type, as each possible combination of the following attributes.
  1. (1)
    SIZE. 7 levels: Indicates if the household includes 1, 2, $\dots$,… , or 7+ members.
  2. (2)
    SPOUSE. 5 levels: Indicates if the householder is not living with a spouse or partner, is married to and living with someone of the same sex, is married to and living with someone of the opposite sex, is cohabiting with a same sex partner, or is cohabiting with an opposite sex partner.
  3. (3)
    RELATIVE. 2 levels: Indicates if a relative of the householder is present in the household or not.
  4. (4)
    OWN_CHILD. 4 levels: Indicates if the householder does not have an own child (either adopted or biological) less than 18 years old present in the household, has an own child less than 6 but does not have an own child who is 6-17, does not have an own child less than 6 but does have an own child who is 6-17, or if the householder has an own child less than 6 years old and also has an own child who is 6-17 years old.
  5. (5)
    CHILD. 4 levels: Indicates if the household does not include a child, includes a child less than 6 but does not include a child who is 6-17, does not include a child less than 6 but does include a child who is 6-17, or if the household includes a child less than 6 years old and also includes a child who is 6-17 years old.
  6. (6)
    MULTIGEN. 2 levels: Indicates if the household is multigenerational or not. A household is considered multigenerational if at least three parent/child generations are present. Households that include either the child and the grandchild of a householder, or include a child and a parent of the householder, are considered multigenerational.
The DHCH schema encodes this information in the single attribute HHTYPE because many of the components that make up HHTYPE are related to one another, and encoding the combination of these variables using a single attribute provides a way to exclude combinations of these variables that are impossible in practice. For example, if the SIZE component of HHTYPE indicates the household is composed of only the householder, then SPOUSE must indicate the householder is not married, RELATIVE must indicate the household does not contain a relative of the householder, OWN_CHILD must indicate the householder does not have an own child, and MULTIGEN must indicate the household is not a multigenerational household. In other words, most elements of the cross product of each of these individual components do not actually correspond to a possible level of HHTYPE. Rather than include all 2,240 elements in this naïve cross product of these components of HHTYPE, we decrease the dimension of the resulting DHCH histogram by only including the 522 valid cross product elements.
The DHCH householder histogram schema is defined by the following attributes and levels.
  1. (1)
    GEOCODE. 5,892,698 levels: Provides the block geographic identification code. [^2] This is the number of blocks in 2020 with at least one housing unit or at least one occupied GQ.
  2. (2)
    SEX. 2 levels: Indicates if the householder is Male or Female.
  3. (3)
    HHAGE. 9 levels: Indicates if the age of the householder is between 15 to 24 years, 25 to 34, 35 to 44, 45 to 54, 55 to 59, 60 to 64, 65 to 74, 75 to 84, or over 85 years.
  4. (4)
    HISPANIC. 2 levels: Indicates if householder is Hispanic/Latino, or not Hispanic/Latino.
  5. (5)
    RACE. 7 levels: Indicates if householder is Black/African American, American Indian/ Alaska Native, Asian, Native Hawaiian/Pacific Islander, White, some other race, or two or more of these race categories.
  6. (6)
    ELDERLY. 4 levels: Indicates if the age of the oldest member of household is less than 60 years, between 60 and 64, between 65 and 74, or over 75 years.
  7. (7)
    TENURE. 3 levels: Indicates if household resides in a housing unit that is owned with a mortgage, owned without a mortgage, or rented.
  8. (8)
    HHTYPE. 522 levels: These levels are described in the previous enumerated list.
In addition to the householder histogram, DHCH DAS implementation generates estimates for the detailed cell counts of the units histogram, which provides counts of housing units and GQs rather than counts of householders. The DHCH units histogram has a schema with only one attribute other than GEOCODE, which is known as TENVACGQ (35 levels). TENVACGQ provides the tenure or vacancy status for each housing unit and the detailed GQ type category for each GQ. Specifically, the first four levels of TENVACGQ correspond to levels of occupied housing units; these levels indicate if a property is mortgaged, owned, rented, or occupied by someone who is not the owner or a renter. The subsequent seven levels correspond to types of vacant housing units; these levels indicate if the housing unit is vacant and for rent, vacant and rented but not occupied, vacant and for sale, vacant and sold but not occupied, vacant because the housing unit is for seasonal, recreational, or occasional use, vacant because the property is for migrant workers, or vacant for other reasons. The subsequent 24 levels correspond to the detailed GQ type categories used in DHCH. The units histogram is estimated at the same time as the householder histogram during the DHCH DAS implementation because a constraint links both of these histograms to one another. This constraint, and all other constraints in the 2020 DHCH production implementation settings, are provided in the list below.
  1. (C1)
    The DHCH MDF is constrained to be consistent with the units redistricting data MDF.
  2. (C2)
    The count of the total number of housing units and number of occupied GQs is held invariant in each block. [^3] Note that the counts of occupied GQs are held invariant in each block, but, for housing units, only the counts of housing units are held invariant in each block. In other words, both the number of vacant and occupied housing units in each block are protected, but their sum is held invariant.
  3. (C3)
    Structural zero constraints are included to ensure consistency of attributes with the SIZE component of HHTYPE. For example, households of size one cannot have a householder with age less than 60 and have a household member greater than 60.
  4. (C4)
    The values of ELDERLY, HHAGE, and HHTYPE are constrained to ensure that the difference between the age of each householder and the implied age of their spouse/partner is no more than 50 years. For example, if the HHAGE attribute indicates the householder is between 15-24 years old and HHTYPE implies that the household contains only the householder and their spouse/partner, then ELDERLY cannot imply that a member of the household is older than 75 years of age.
  5. (C5)
    The value of ELDERLY is constrained to be at least as high as HHAGE for each record.
  6. (C6)
    Age differences between parents and children are constrained to be less than the upper bound on these differences that is used in the CEF edit constraints. For example, these constraints rule out households with a householder who is over 75 years old and that includes a child of the householder who is less than 6.444The CEF edit constraints bound the maximum age difference between a mother and a child to 50 years and the maximum age difference between a father and a child to 69 years.
  7. (C7)
    The DHCH householder and unit histograms are constrained to be consistent with one another; the count of householders who own their home with a mortgage, own without a mortgage, and do not own their home in the householder histogram are each constrained to be equal to the count of homes that are owned with a mortgage, owned without a mortgage, and that are not owned by the current occupants in the units histogram, respectively.
2.2. DHCP Schema
For the persons data universe, the DHCP schema is defined by the following attributes and levels.
  1. (1)
    GEOCODE. 5,892,698 levels: Provides the block geographic identification code.
  2. (2)
    RELGQ. 42 levels: For respondents in households, the relationship to householder is encoded by the first 18 levels, and, for the remaining respondents, the detailed GQ type is encoded by the final 24 levels.
  3. (3)
    SEX. 2 levels: Indicates if the respondent is Male or Female.
  4. (4)
    AGE. 116 levels: Contains one level for each single year of age between 0 and 115.
  5. (5)
    HISPANIC. 2 levels: Indicates if the respondent is Hispanic/Latino, or not Hispanic/Latino.
  6. (6)
    CENRACE. 63 levels: Includes every combination of Black/African American, American Indian/Native Alaskan, Asian, Native Hawaiian/Pacific Islander, White, and some other race, except “none of the above.”
The invariant constraints for the 2020 DHCP DAS implementation are described in the following list.
  1. (Inv1)
    The DHCP MDF is constrained to be consistent with the persons redistricting data, e.g., any tabulation in the redistricting data file will be the same as this tabulation derived using the DHCP MDF.
  2. (Inv2)
    The count of occupied GQs for each detailed GQ type in each block is held invariant. In other words, for a given block and a given detailed GQ type, we require at least the same number of records in this detailed GQ type and block as the number of occupied GQs of this type in the block. In addition, the number of records in each GQ is constrained to be at most 99,999, and the number of records in each housing unit is also constrained to be at most 99,999.
  3. (Inv3)
    Depending on the value of RELGQ of each record, the AGE attribute of the record is constrained to satisfy the lower and upper bounds given in Table 1, which are also edit constraints that are used when creating the CEF.
Relationship to Householder or GQ Type
Age Range
Householder
15-115
Spouse or unmarried partner of householder
15-115
Adopted or Biological child
0-89
Stepson/stepdaughter
0-89
Brother or sister
0-115
Parent
30-115
Grandchild
0-74
Parent-in-law
30-115
Son-in-law/daughter-in-law
15-89
Other relative
0-115
Housemate/roommate
0-115
Foster child
0-20
Other non-relative
0-115
Non-military correctional facilities (GQ Types 101-105)
15-115
Group home/correctional/treatment facilities for juveniles (GQ Types 201-203)
0-25
Nursing facilities (GQ Type 301)
20-115
Hospitals/hospices (GQ Types 401-403)
0-115
Residential schools for people with disabilities (GQ Type 405)
3-30
College/university student housing (GQ Type 501)
16-65
Military Housing (GQ Types 106, 404, 601, 602)
17-65
Adult group homes/treatment centers (GQ Types 801, 802)
16-115
Maritime/merchant vessels (GQ Types 900)
16-75
Other non-institutional (GQ Types 701-706, 901-904)
0-115
Table 1. This table provides the feasible age ranges by relationship to householder and GQ type.
2.3. Types of Constraints Within the DAS
For both the DHCH and DHCP DAS implementations, it is worthwhile to separate constraints into two categories. First, edit and structural zero constraints are imposed on the output of the DAS to ensure that the constraints imposed during the creation of the CEF are also satisfied in the output MDF. These constraints can be viewed as the requirements provided by subject matter experts for the face validity of the microdata. These constraints are also data-independent and publicly available. As a result, and also since our privacy guarantees are specified with respect to the records in the CEF, these constraints do not limit the types of inferences on the CEF records that we protect against. Some examples of constraints in this category include that each householder record in the DHCP MDF is constrained to be at least 15 years of age, and, in the DHCH MDF, that each household of size one with a householder with age less than 60 is constrained to have ELDERLY indicate that no household member is over 60 years of age. Other examples of these constraints include constraints (C3-C7) and (Inv3) in the list of DHCH and DHCP constraints above, respectively.
Second, our production DHC implementations also impose several constraints, which we call invariant constraints, that are dependent on the input data files of the DAS DHC implementations rather than on publicly-available data. Specifically, the invariant constraints imposed on the production DHC implementations are (A) the total populations of each state, Washington D.C., and Puerto Rico, (B) the tabulation block locations of each occupied GQ by detailed GQ type, e.g., the location of each nursing facility, and (C) the tabulation block locations of each housing unit. The first set of constraints is imposed in the persons universe of the production 2020 Redistricting Data File DAS implementations. This is carried forward into the production DHCP implementation through our constraints that ensure the DHC outputs are consistent with the redistricting data file. While we protect inferences on whether a given a housing unit is occupied or vacant, we do not provide this same protection to GQs because unoccupied GQs are out of scope for the 2020 Census and the DAS has no access to that information. More detail on the rationale for the inclusion of these invariant constraints is provided by [14].
It is possible to use these set of invariant constraints to formulate inferences on individuals that are not protected by the DAS. For example, if a privacy attacker knows an individual resides in a given block that does not have an occupied nursing facility in the DHCH units table, it is possible for the privacy attacker to infer that the individual is not in a nursing facility.
2.3.1. Unused Constraints in DHC
Note that the constraints imposed within the DHC DAS implementation do not avoid all inconsistent counts. For example, since the persons and units universe are processed separately from one another, there are some blocks with occupied housing units in the DHCH MDF but without any population in the DHCP MDF. Inconsistent counts such as these are often a feature of sparsely populated areas, as these areas often pose a higher risk of disclosure. The presence of inconsistent counts is a warning to the user about the presence of disclosure avoidance. It is worth noting that inconsistent counts are an accepted feature in other countries, such as in the UK Census [21, 24].
2.3.2. How Constraints are Imposed
As described previously, the primary way that fitted histograms are computed to satisfy the DHCH and DHCP constraints is through a series of least squares and mixed integer programming optimization solves. These optimization problems are described in detail by [14]. As is also described in [14], we require the constraint matrix for the MIP optimization problems to satisfy a condition known as total unimodularity. This restriction ensures these MIP solves are computationally tractable; for more detail on total unimodularity and the sufficient condition we use to ensure it, see [14, 15].
With the exception of the DHCP constraints described in (Inv2) from Section 2.2, i.e., the DHCP upper and lower GQ count constraints, all other DHCH and DHCP constraints are imposed using these optimization solves. A different approach is required for DHCP constraints described in (Inv2) because the sufficient condition we use to ensure total unimodularity would not hold if these constraints were added to these MIP optimization problems. For this reason, we created a post-processing method called the DHCP optimization heuristic that improves the feasibility for the DHCP constraint (Inv2) of the histogram estimate that is output from the least squares and MIP optimization solves. The DHCP optimization heuristic works by iteratively choosing pairs of histogram elements within the same histogram of a single geographic unit and then considering increasing one count by one and decreasing the other by one. If changing the histogram counts in this way improves feasibility of a DHCP constraint in (Inv2) and does not result in any of the remaining constraints being infeasible, this change to these two histogram counts is accepted, and otherwise the change is rejected and no change is made to these two histogram counts. Afterward, the algorithm goes on to consider the next change in a pair of histogram counts.
In more granular geographic units, i.e., optimized block groups and census blocks, the DHCP optimization heuristic chooses pairs of histogram cells for which all attributes other than RELGQ are identical. Since it is possible for one of the DHCP constraints described in Section 2.2 to become infeasible for all such changes to pairs of histogram counts, the DHCP constraint described in (Inv2) does not always hold for each of these geographic units, as described in the next example.
Example 1.
Suppose the census block histogram estimate that is input into the DHCP optimization heuristic encodes one person who is 31 years old that lives in a hospital, hospice, or treatment facility (GQ types 401-404), but no people residing in residential schools for people with disabilities (GQ type 405). Also suppose the DHCP upper and lower GQ count constraints described in (Inv2) require that at least one respondent resides in GQs of type 405 and that no respondents reside in GQs of types 401-404. In order to maintain feasibility for the DHCP constraints described in (Inv1), i.e., the constraints that ensure consistency with the redistricting data file MDF, we can only consider switching the respondent who currently has a GQ type of 401-404 to a GQ type of 405, since the number of respondents estimated to be in GQ types 401-405 is held fixed by this DHCP constraint. However, switching the GQ type of this respondent from a GQ type in the range 401-404 to 405 would impair the feasibility of the DHCP constraints descried in (Inv3) because this constraint requires that respondents with a GQ type of 405 be less than 30 years old. Since all changes that we consider for this block geographic unit result in the feasibility of either the DHCP constraint described in (Inv1) or in (Inv3) to fail, the DHCP optimization heuristic would not make any changes to the input histogram estimate in this case.
However, in the case of all geographic units that are less granular than optimized block groups, we also consider switching respondents between pairs of histogram cells with different values for the attribute AGE. As described in the next example, this results in all of the DHCP constraints described in (Inv1-Inv3) being satisfied for each geounit.
Example 2.
We consider the same setting as Example 1, but, rather than considering the case in which the geographic unit is a census block, we suppose it is a state geounit instead. In this case the DHCP optimization heuristic considers changes to both the RELGQ attribute and the AGE attribute. As described in Example 1, only changes to RELGQ that result in the respondent’s GQ type being changed to GQ type 405 are considered. Since one attribute in the redistricting data file is VOTING_AGE, which is one if the respondent is at least 18 years old and zero otherwise, the DHCP constraints described in (Inv1) require that only changes to the AGE attribute that change the age of the respondent from 31 to any age of at least 18 are considered. In the case of this state geographic unit, the DHCP optimization heuristic considers pairs of histogram counts until it eventually accepts a change that results in this record’s GQ type being switched to 405 and its age being switched to a value between 18 and 30.
3. Computational Requirements for the DHC DAS
[14] provide detail on the computational environment used by the redistricting data file DAS implementation, including a list of software dependencies, AWS cluster configurations, and Spark settings. In this section, we describe the computational environment used for our DHC DAS implementations, with a particular focus on the aspects that we changed after publication of [14].
The factors that did not change include the software dependencies, i.e., our production code base is written in Python version 3.7, we use Gurobi 9.1, and the system ran on Amazon Web Services’ (AWS) using Elastic Map Reduce (EMR) version 6.2 and Apache Spark version 3.0.1. The DAS is run on EMR clusters built from AWS nodes with 96 cores and 768 GiB of RAM. The DHC executions were carried out on clusters with 20 (core and task) worker nodes and one EMR master node.
Several changes were also made to avoid computational issues due to the DHC schemas being much larger than their redistricting data counterparts. First, due to the memory requirements of the large optimization solves used in DHC executions, we increased executor memory and overhead memory to 423 GiB and 330 GiB, respectively. Our typical DHC executions used 20 executors, 40 cores per executor, and 40 GiB of driver memory. We also made several changes to our cluster configuration settings, including increasing the heap sizes of the yarn node manager and resource manager to 8 GiB and 4 GiB, respectively, increasing the EBS storage to 16 TiB, and switching to a G1 garbage collection method.
4. DHC Accuracy Experiments and PLB Allocations
In preparation for the 2020 DHC production DAS implementations, the DAS team performed a series of accuracy experiments using 2010 Census inputs in order to support recommendations for the Census Bureau’s final production settings of the 2020 DHC production implementations, including the PLB allocations. The DHC accuracy experiments were similar to those used to support the final production settings for the 2020 Redistricting Data (P.L. 94-171) Summary File production DAS implementations [14]. Specifically, these accuracy experiments used an iterative process consisting of DAS configuration updates based on input from subject-matter experts (SMEs) on the accuracy of tabulations from prior experiment iterations. The Census Bureau internal SMEs included those from the Population Division, and Social, Economic, and Housing Statistics Division (SEHSD).
External SMEs also provided feedback based on the output of two DHC demonstration data products (DDPs) based on 2010 Census input data released on March 16, 2022 and August 25, 2022, respectively. [^5] Despite the name “demonstration data products,” these releases were not meant to demonstrate the final production settings we intended to use. These were instead intended to demonstrate performance of the DHC DAS implementation for the settings we were currently testing at the time in our DHC accuracy experiments. This includes feedback provided through a National Academies of Science, Engineering, and Medicine’s Committee on National Statistics (CNSTAT) meeting in June of 2022 [18], academic conferences, and external stakeholders independently contacting the Census Bureau regarding the accuracy of these 2010 DDPs for specific use cases. In addition to these two DDPs, the DAS output based on 2010 Census input data and using the 2020 DHC production implementation settings was also released on April 3, 2023. All of these releases are currently available at [26].
The result of these accuracy experiments was the 2020 DHC production configuration parameters, including strategy matrices for both DHCH and DHCP, which refer to the marginals that are chosen to become noisy measurements as well as the PLB allocation of each marginal that defines the scale of the noise added to the marginal for each geographic level. As described in [14] in more detail, we use the $\rho-$zero-concentrated differential privacy ($\rho-$zCDP) accounting framework to specify these PLB values [16]. Each strategy matrix described in the next two subsections provides the initial PLB allocations for each geographic unit on the spine, prior to the spine optimization routines’ alteration of these allocations. Specifically, as described by [19] in more detail, in the case of parent and child geographic units for which the parent geographic unit has only one child, the spine optimization routines alter the initial PLB allocations of these geographic units. Aside from dictating the scale of the noise for each noisy measurement used within the DAS, the PLB allocations are also useful because they can be used to derive upper bounds on the power of certain hypothesis tests on any given individual record in the CEF; for more detail, see [23].
4.1. DHCH Production PLB Allocations
The marginal query groups used in the DHCH strategies are provided in Table 2. Many of these marginals are defined by the attributes in the Section 2.1. The definitions of the remaining marginals, i.e., recodes, are given in the list below.
  1. (1)
    TENURE_2LEVELS. 2 levels: Indicates if an occupied housing unit is owned or rented.
  2. (2)
    PARTNER_TYPE_CHILD_STAT. 12 levels: This HHTYPE attribute marginal can be viewed as the cartesian product of two variables. The first variable contains three levels that indicate if a householder has no spouse or partner, is married, or is cohabiting with a partner. The second variable is given by the HHTYPE attribute OWN_CHILD, as defined in Section 2.1.
  3. (3)
    HHTYPE_RECODE. 183 levels: Defined as the cross product of HHTYPE attributes, MULTIGEN $\times$×SPOUSE×$\times$×SIZE×$\times$×OWN_CHILD.
Table 2. Production DHCH Strategy Queries
Query ID
Marginal Query Group Name
Cells
1H
MULTIGEN×$\times$×HISPANIC×$\times$×TENURE_2LEVELS
8
2H
HISPANIC×$\times$×TENURE_2LEVELS
4
3H
HISPANIC×$\times$×TENURE_2LEVELS×$\times$×RACE
28
4H
PARTNER_TYPE_CHILD_STAT×$\times$×SEX×$\times$×TENURE_2LEVELS
48
5H
SPOUSE×$\times$×HISPANIC×$\times$×TENURE_2LEVELS
20
6H
SEX×$\times$×HISPANIC×$\times$×TENURE×$\times$×RACE×$\times$×HHTYPE_RECODE
15,372
7H
SEX×$\times$×HISPANIC×$\times$×TENURE×$\times$×RACE×$\times$×HHAGE×$\times$×HHTYPE_RECODE
1,383,48
8H
TENURE×$\times$×HHAGE×$\times$×HHTYPE_RECODE
4,941
9H
SEX×$\times$×HISPANIC×$\times$×TENURE×$\times$×RACE×$\times$×ELDERLY×$\times$×HHAGE×$\times$×HHTYPE
1,578,528
10H
TENVACGQ
35
Each marginal query group in Table 2 is assigned a query ID, and the PLB allocations for the production US DHCH implementation for each query ID and each geographic level are given in Table 3. The values in the table below are the actual $\rho$ allocations multiplied by 10,000; in other words, the actual $\rho$ allocations used in the DHCH DAS production settings are given by the values in the table below divided by 10,000. The total value of $\rho$ allocated to all marginal query groups for our production DHCH settings was 15401/2000 ($\approx 7.70$).
Table 3. Production US DHCH $\rho$ Allocations ×10,000$\times$ 10,000× 10 , 000 (i.e., divide by 10,000 to obtain the $\rho$ value)
Query ID
US
State
County
Prim
TSG
TS
OBG
Block
1H
792
714
986
973
973
1,150
291
11
2H
212
193
266
214
214
310
78
3
3H
212
191
1,500
1,500
1,500
1,750
442
17
4H
792
1428
1,972
1,946
1,946
2,300
581
22
5H
792
714
3,944
973
973
4,601
1,162
44
6H
257
3,216
1,380
973
973
1,610
407
15
7H
257
3,216
1,380
973
973
1,610
407
15
8H
212
193
266
214
214
310
78
3
9H
1,049
3,931
1,380
973
973
1,610
407
15
10H
259
3,216
1,380
973
973
1,610
407
15
*The prim, tract subset group (TSG), tract subset (TS), and optimized block group (OBG) geographic levels are not in the tabulation geographic spine and are only used internally in the TDA.
4.2. DHCP Production PLB Allocations
As described in the previous section for the DHCH schema, the marginal query groups used in the DHCP strategies are provided in Table 4. Many of these marginals are defined by the attributes in the Section 2.2. The definitions of the remaining marginals, i.e., recodes, are given in the list below.
  1. (1)
    AGE_18_64. 3 levels: Indicates which of the following categories the respondent’s age is in AGE $\leq 17,$ $18\leq$ AGE $\leq 64,$ or AGE $\geq 65.$
  2. (2)
    RELGQ_4_GROUPS. 4 levels: Indicates if the respondent is either a householder, a non-householder member of a household, a resident of an institutional GQ, or a resident of a non-institutional GQ.
  3. (3)
    AGE_38_GROUPS. 38 levels: Indicates which of the following categories the respondent’s age is in AGE $=0,$ AGE $=1,\dots,$ AGE $=14,$ AGE $\in[15,17],$ AGE $\in[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],\dots,$ AGE $\in[55,59],$ AGE $\in[60,61],$ AGE $\in[62,64],$ AGE $\in[65,66],$ AGE $\in[67,69],$ AGE $\in[70,74],$ AGE $\in[75,79],\dots,$ or AGE $\geq 110.$
  4. (4)
    GQ_CONSTR_GROUPS. 6 levels: Indicates if the respondent is a resident of a GQ with a lower-bound constraint on AGE given by 0, 3, 15, 16, 17 years of age, as described in Table 1, or if the respondent is a member of a household. In other words, GQ_CONSTR_GROUPS indicates if the respondent’s GQ type is in $\{201,202,203,301,401,402,403,701\},\{405\},$$\{101,102,103,104,105\},$$\{501,801,802,900,901,702,704,706,903,904\},\{106,404,601,602\},$ or if the respondent is in a household.
  5. (5)
    AGE_10_GROUPS. 10 levels: Indicates which of the following categories the respondent’s age is in AGE $\leq 3,$ AGE $=[3,14],$ AGE $=15,$ AGE $=16,$ AGE $=17,$ AGE $=[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],$ AGE $\in[30,34],$ or AGE $\geq 35.$
  6. (6)
    POP_SEHSD_TARGETS_RELSHIP. 19 levels: Indicates if the respondent is in a GQ or, for residents who are members of a household, which of the 18 household RELGQ categories the respondent belongs to.
  7. (7)
    AGE_40_GROUPS. 40 levels: Indicates which of the following categories the respondent’s age is in AGE $=0,$ AGE $=1,\dots,$ AGE $=17,$ AGE $\in[18,19],$ AGE $\in[20,24],$ AGE $\in[25,29],\dots,$ AGE $\in[55,59],$ AGE $\in[60,61],$ AGE $\in[62,64],$ AGE $\in[65,66],$ AGE $\in[67,69],$ AGE $\in[70,74],$ AGE $\in[75,79],\dots,$ AGE $\in[105,109],$ or AGE $\geq 110.$
  8. (8)
    RELSHIP_GQ. 25 levels: For respondents in GQs, indicates the respondent’s major GQ type, and, for respondents who are members of households, indicates the respondent’s household RELGQ category.
Table 4. Production DHCP Strategy Queries
Query ID
Marginal Query Group Name
Cells
1P
AGE_18_64$\times$×RELGQ_4_GROUPS
12
2P
AGE_18_64$\times$×SEX
6
3P
AGE_38_GROUPS×$\times$×SEX
76
4P
HISPANIC×$\times$×SEX
4
5P
SEX×$\times$×RELGQ_4_GROUPS
8
6P
GQ_CONSTR_GROUPS×$\times$×AGE_10_GROUPS
60
7P
POP_SEHSD_TARGETS_RELSHIP
19
8P
HISPANIC×$\times$×SEX×$\times$×AGE_40_GROUPS×$\times$×RELSHIP_GQ×$\times$×CENRACE
252,000
9P
RELGQ×$\times$×AGE_40_GROUPS×$\times$×HISPANIC×$\times$×CENRACE×$\times$×SEX
423,360
10P
RELGQ×$\times$×SEX×$\times$×AGE×$\times$×HISPANIC×$\times$×CENRACE
1,227,744
The PLB allocations for the production US DHCP implementation for each query ID in Table 4 and each geographic level are given in Table 5. As is done in the previous subsection, the values in the table below are the actual $\rho$ allocations multiplied by 10,000; in other words, the actual $\rho$ allocations used in the DHCP DAS production settings are given by the values in the table below divided by 10,000. The total value of $\rho$ allocated to all marginal query groups for our production DHCP settings was 24811/5000 ($\approx 4.96$).
Table 5. Production US DHCP $\rho$ Allocations ×10,000$\times$ 10,000× 10 , 000
Query ID
US
State
County
Prim
TSG
TS
OBG
Block
1P
73
999
310
478
478
868
430
11
2P
73
999
310
478
478
868
430
11
3P
73
999
310
478
478
868
430
11
4P
73
999
310
478
478
868
430
11
5P
73
999
310
478
478
868
430
11
6P
292
3,996
1,240
1,912
1,912
3,472
430
11
7P
73
999
1,240
478
478
3,472
430
11
8P
73
999
310
478
478
868
430
11
9P
73
999
310
478
478
868
430
11
10P
73
999
310
478
478
868
430
11
*The prim, tract subset group (TSG), tract subset (TS), and optimized block group (OBG) geographic levels are not in the tabulation geographic spine and are only used internally in the TDA.
5. Summary Metrics for 2020 DHC Production Implementations
[25] provides data users with $\rho-$zCDP summary metrics for the 2020 DHC release using a discrete Gaussian mechanism. The total PLB for all the summary metrics in this release was $\rho=0.7.$ We used a formally private method to produce these metrics because they summarize errors of the 2020 DHC release MDF-based tabulations with respect to CEF-based tabulations. This section provides the approximate mean absolute error (MAE) $\rho-$zCDP metrics from this document for several important DHC tabulations. The MAE of a query in a given geographic level is defined by first finding the query error in each geographic unit in the geographic level; afterward, the MAE is defined as the arithmetic mean of the absolute value of these errors. The PLB allocated to each of the state geographic level MAEs was set so that the reported approximate MAE was within 0.5 units of the unprotected CEF-based MAE value with probability 0.9. Likewise, the PLB allocated to the remaining summary metrics was set so that the reported approximate MAE was within 0.15 units of the unprotected CEF-based MAE value with probability 0.9. Many other $\rho-$zCDP approximate summary metrics in addition to the ones that are provided here are available in [25].
Tables 6 through 7 provide the MAE values of tabulations in the DHCH universe. As described previously, counts in this universe can be viewed as counts of households; alternatively, since there is one householder per household, counts in this universe can also be viewed as counts of householders. For the DHCP universe, Tables 8 and 9 provide the MAE of counts of individuals grouped by relationship to the householder and by their age group, respectively.
Table 6. Production US 2020 DHCH: Presence of Own Child of Householder MAE
Group
States
Counties
Places
Tracts
Own children under 6 years only
10.39
5.49
2.99
4.10
Own children between 6 and 17 years only
10.98
5.71
3.30
4.59
Own chilren under 6 and between 6 and 17 years
14.88
5.71
3.11
4.09
No own children under 18 years
12.84
6.05
4.03
5.20
Table 7. Production US 2020 DHCH: Tenure by Race MAE
Group
States
Counties
Places
Tracts
Owner occupied
White alone
22.78
5.56
2.70
3.27
Black/African American alone
27.76
4.34
1.54
2.30
American Indian/Alaska Native alone
28.88
3.40
1.01
1.50
Asian alone
19.39
3.33
1.19
1.99
Native Hawaiian/OPI* alone
15.86
1.75
0.38
0.47
Some Other Race alone
13.33
3.81
1.60
2.24
Two or More Races
15.08
4.20
2.12
2.74
Renter occupied
White alone
17.08
4.70
2.26
2.95
Black/African American alone
13.92
3.94
1.42
2.26
American Indian/Alaska Native alone
20.80
2.83
0.99
1.47
Asian alone
12.10
2.67
0.95
1.63
Native Hawaiian/OPI* alone
13.47
1.59
0.42
0.61
Some Other Race alone
11.82
3.29
1.41
2.15
Two or More Races
12.24
3.77
1.66
2.49
*OPI stands for Other Pacific Islander.
Table 8. Production US 2020 DHCP: Relationship to Householder MAE
Group
States
Counties
Places
Tracts
Householder
23.02
6.70
4.32
4.26
Opposite-sex husband/wife/spouse
8.43
4.50
3.39
3.26
Opposite-sex unmarried partner
10.06
3.45
2.40
2.37
Same-sex husband/wife/spouse
8.39
2.58
1.44
1.47
Same-sex unmarried partner
6.92
2.33
1.37
1.35
Biological son or daughter
18.67
5.92
3.75
3.68
Adopted son or daughter
9.98
3.64
2.17
1.98
Stepson or stepdaughter
9.59
3.75
2.29
2.17
Brother or sister
8.35
3.30
2.15
2.18
Father or mother
7.02
3.08
2.08
2.06
Grandchild
11.35
3.91
2.60
2.58
Parent-in-law
6.25
2.63
1.47
1.57
Son-in-law or daughter-in-law
5.31
2.70
1.63
1.67
Other relative
11.02
3.93
2.33
2.33
Roommate or housemate
9.73
3.80
2.15
2.30
Foster child
6.61
2.58
1.46
1.23
Other nonrelative
10.80
3.84
2.37
2.21
Table 9. Production US 2020 DHCP: 5 Year Age Groups MAE
Group
Counties
Places
Tracts
AIR/ORT*
Under 5 years
9.52
4.47
5.28
3.95
5 to 9 years
10.64
4.36
5.84
4.35
10 to 14 years
10.63
4.50
5.95
4.61
15 to 19 years
8.31
4.52
4.39
2.84
20 to 24 years
4.57
2.78
3.06
2.08
25 to 29 years
4.59
2.80
3.09
2.17
30 to 34 years
4.39
2.77
3.02
2.12
35 to 39 years
6.56
3.01
4.06
3.29
40 to 44 years
6.36
2.94
4.00
2.86
45 to 49 years
6.25
2.93
3.95
2.89
50 to 54 years
6.22
2.99
3.96
3.20
55 to 59 years
6.17
3.07
3.96
3.35
60 to 64 years
8.08
3.80
4.99
3.33
65 to 69 years
7.90
3.63
4.75
3.53
70 to 74 years
6.01
2.84
3.67
2.83
75 to 79 years
5.87
2.71
3.50
2.76
80 to 84 years
5.88
2.55
3.33
2.54
85 to 89 years
5.39
2.29
3.10
1.90
90 to 94 years
5.29
2.19
2.75
1.55
95 to 99 years
4.77
1.69
2.00
0.99
100 to 104 years
3.14
0.94
0.77
0.38
105 to 109 years
1.29
0.27
0.14
0.40
over 109 years
0.64
0.03
0.05
0.19
*AIR/ORT stands for Federal American Indian Reservations/Off-Reservation Trust Lands.
6. Conclusion and Progress after the 2020 Census DHC Release
As described in Section 1, one important advantage of formally private mechanisms like the DAS is that they are sufficiently transparent to allow users to construct confidence intervals about the private CEF-based tabulations in a straightforward manner. To describe how this can be done, more background on how the strategy matrices described in the last section are used within the DAS is helpful. Specifically, after spine optimization, the noisy measurements for each tabulation and each geographic unit are defined by adding discrete Gaussian noise to the CEF-based tabulation of the given geographic unit. If the PLB allocation of the geographic unit and tabulation pair is given by $\rho,$ the mean-zero noise added to each level of the CEF-based tabulation is given by a realization of a random variable from the discrete Gaussian distribution, $N_{\mathbb{Z}}(\mu=0,\sigma^{2}=1/\rho),$ which is defined so that its probability mass function is proportional to $f(x)=\exp(-x^{2}/2\sigma^{2})$ for each $x\in\mathbb{Z}$ and zero otherwise [17]. We call the collection of all such noisy measurements for a given DAS implementation the noisy measurement files (NMFs). The NMFs provide sufficient information to construct the likelihood function for each geographic unit and each tabulation for which a noisy measurement is observed. This in turn can be used to generate point estimates and confidence intervals via simulation; alternatively, the underlying discrete Gaussian random variables can be modeled using continuous Gaussian random variables to derive closed form confidence intervals; see for example [20].
To support research and development on tools that would provide users inferences on underlying CEF-based tabulations, the Census Bureau has released NMFs for both the final 2010-CEF-based DDP production settings DAS implementations and the 2020 production DAS implementations as experimental data products. [26] provides instructions for downloading the NMFs of both of these implementations. Note that this method of producing confidence intervals using the NMFs typically does not result in an interval that is centered around the count implied by the MDF output from the DAS. Instead of relying on tabulations of counts from the MDF that is output from the DAS for the center of these confidence intervals, these intervals are centered around the inverse-variance-weighted mean of values derived from the NMFs, which, unlike MDF-based tabulations, provides an unbiased estimate of the count derived from the CEF.
The DAS team is also developing methods that provide confidence intervals for tabulations derived using the output of the production Redistricting and DHC MDFs. These confidence intervals are centered around the tabulations derived using the production MDFs, so they also provide a measure of the uncertainty due to disclosure avoidance of the published DHC tabulations.